[figure: 9.7 cm]

Phase transitions and crossovers in anisotropic two-dimensional

quantum antiferromagnets

T. Roscilde^{a},A. Cuccoli^{b}, V. Tognetti^{b}, P. Verrucchi^{b}, R. Vaia^{c} aDipartimento di Fisica ”A. Volta” dell’Università di Pavia and INFM, via A. Bassi 6,

I-27100 Pavia, Italy bDipartimento di Fisica dell’Università di Firenze and INFM, via G. Sansone 2, I-50019 Sesto Fiorentino, Italy c Istituto di Elettronica Quantistica del Consiglio Nazionale delle Ricerche, via Panciatichi 56/30, I-50127 Firenze, and INFM

We study the thermodynamic properties of anisotropic two-dimensional quantum antiferromagnets, of both easy-plane and easy-axis nature, as modeled by the XXZ model in zero field. We make use of the Quantum Monte Carlo method based on the continuoustime loop algorithm. The quantum simulations, performed in the limit of very small anisotropy and in the case S = 1/2, reveal that both the Ising and Berezinskii-Kosterlitz-Thouless universality class (for the easy-axis and easy-plane models, respectively) are well observed and the critical temperature remains finite even for anisotropies as small as 10^{−3}, which is the typical order magnitude for the anisotropies measured in the real layered compounds; this result rules out the possibility of quantum disorder to appear at finite temperature before the isotropic limit is reached. The phase diagram of the S = 1/2 2d-XXZ antiferromagnet in the region of small anisotropies is then presented. Finally, in the easy-plane case direct comparison is made with experimental data of the S = 1/2 layered antiferromagnet Sr_{2}CuO_{2}Cl_{2}, and evidence is given for a Heisenbergto-XY crossover in this compound, as conjectured by previous works.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : oral

Corresponding author : Tommaso Roscilde

Address for correspondence : Tommaso Roscilde c/o Paola Verrucchi Dipartimento di Fisica dell’Università di Firenze via G. Sansone 1, I-50019 Firenze, Italy

Email address : roscilde@fi.infn.it